OpenStudy (anonymous):

Solve for x: 2 over 5 (x − 4) = 2x.

OpenStudy (anonymous):

\[\frac{2}{5(x-4)}=2x\] multiply both sides by 5(x-4) \[\frac{2}{5(x-4)}*5(x-4)=2x*5(x-4)\]

OpenStudy (anonymous):

simplify 2=10x(x - 4)

OpenStudy (anonymous):

thx

OpenStudy (anonymous):

will you continue on your own?

OpenStudy (anonymous):

i forgot how to simplify

OpenStudy (anonymous):

how do you forget how to simplify

OpenStudy (anonymous):

idk i forgot

OpenStudy (anonymous):

can someone please help?

OpenStudy (campbell_st):

by simplifying all you need to do is ditribute the 10x so \[2 = 10x^2 -40x\] rewrite the equation and you get \[10x^2 - 40x - 2 = 0\] so you can now solve for x, using the general quadratic formula so make it a little easier, divide every term by 2 \[5x^2 - 20x - 1 = 0\]

OpenStudy (anonymous):

i got 13

OpenStudy (anonymous):

am i wrong?

OpenStudy (campbell_st):

well given its a quadratic equation, you need 2 answers... and you need the formula \[x = \frac{-b \pm \sqrt{b^2 - 4ac}}{2a}\] you have a = 5, b = -20 and c = -1